OpenStudy (chrispaul03):

WILL FAN AND MEDAL x/4=5

OpenStudy (anonymous):

Yup, whatcha need help with!

OpenStudy (chrispaul03):

x/4=5

OpenStudy (anonymous):

oh, i see! g

OpenStudy (anonymous):

you have to simplify it?

OpenStudy (chrispaul03):

how do is simplify it

OpenStudy (kendricklamar2014):

Is the question: \[\frac{ x }{ 4 } = 5\]

563blackghost (563blackghost):

Does it look like this? \(\huge{\frac{x}{4}=5}\) If so the way we would solve this is that we would multiply \(\large{4}\) to get \(\large{x}\) by itself.... \(\huge\color{purple}{\frac{x}{4} \times 4 = 5 \times 4}\) \(\huge\color{green}{~~~~~~~~~~ \downarrow}\) \(\huge\color{red}{~~~~~~x~=~?}\)